# Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution

### v4

@@ Abstract
This study proposes a theory of unsupervised super-resolution data assimilation (SRDA) using conditional variational autoencoders (CVAEs). The theory is based on an evidence lower bound that leads to an objective function for unsupervised learning. This function can be reduced to the objective function of a traditional data assimilation method, namely the three-dimensional variational (3D-Var) formalism. Similar to 3D-Var, it is essential to use a non-diagonal background error covariance matrix to assimilate distant observations. Instead of using such a non-diagonal matrix, we extend the proposed theory by using the non-locality of super-resolution (SR). For linear SR, SR operators serve as background error covariance matrices. For nonlinear SR, error backpropagation through SR neural networks induces covariance structures in inference. SRDA can be performed using CVAEs because the loss function for CVAEs is generally an evidence lower bound. By incorporating the SR neural network into the CVAE, the encoder estimates the high-resolution (HR) analysis from HR observations and low-resolution (LR) forecasts. The decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis. The effectiveness of SRDA was evaluated through numerical experiments using an idealized barotropic ocean jet system. Compared to inference with an ensemble Kalman filter, SRDA demonstrated superior accuracy in HR inference. SRDA was also computationally efficient because it does not require HR numerical integration or ensemble calculations. The findings of this study provide a theoretical basis for integrating SR and DA, which will stimulate further research in this direction.
keywords:
super-resolution , data assimilation , unsupervised learning , evidence lower bound , background error covariance , variational autoencoder
journal: Comput. Methods Appl. Mech. Engrg.corresponding_author_infocorresponding_author_infofootnotetext: Corresponding author. 2-12-1-i9-111 Ookayama, Meguro-ku, Tokyo, 1528550, Japan. Tel.: +81 3 57343173 (office). Email address: yasuda.yuki@scrc.iir.isct.ac.jp
[…]
[inst1] organization=Supercomputing Research Center, Institute of Integrated Research, Institute of Science Tokyo, addressline=2-12-1 Ookayama, Meguro-ku, city=Tokyo, postcode=1528550, country=Japan
1 Introduction
@@
Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD) [1, 2]. Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time. Uncertainties in model parameters may also lead to discrepancies from reality. DA reduces these errors by correcting forecast values or model parameters based on observational data, thereby being essential for accurate predictions, such as numerical weather prediction.
@@
Recent advances in deep learning are impacting various scientific fields and are also found in DA. One reason is that both DA and deep learning use backpropagation based on adjoint equations [3, 4, 5]. Another reason is the use of the Bayesian approach [2, 5, 6]. This approach models the true states of fluid systems by probability distributions based on known data, such as observations and fluid model outputs.
@@
A typical neural network using the Bayesian approach is the variational autoencoder (VAE) [7, 8, 9]. Like standard autoencoders [10], a VAE consists of an encoder and a decoder, both of which are implemented by neural networks. Typically, the encoder transforms the input data into a compressed representation in the latent space, while the decoder reconstructs the original data from these latent variables. A major extension of VAEs is the conditional VAE (CVAE), which is based on conditional probability modeling [11, 12]. This conditioning allows us to model target variables, such as the true states of systems, conditioned on known data, such as observations.
@@
Applications of VAEs in DA can be categorized into two groups. The first group employs VAEs for dimensionality reduction [13, 14, 15]. In this approach, the encoder maps fluid model states into a low-dimensional latent space. Statistical DA methods, such as Kalman filters, are then applied in this space. Due to the reduced dimensionality, the computational cost for DA is significantly lowered. This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks [16, 17, 18, 19, 20, 21]. The second group employs VAEs as generators that learn the probability distributions of flow fields [22, 23, 24]. In this approach, a fluid model state is encoded, and noise is then added in the latent space, creating various potential latent variables. Ensemble members are obtained by decoding these latent variables back into the physical space. This process facilitates fast computation of covariance matrices used in statistical DA methods. A common characteristic in both groups is that VAEs do not directly perform DA; rather, VAEs are used to make statistical DA methods more efficient. We proposes a new method, categorized into the third group, where VAEs directly perform DA.
@@
In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state [1, 2]. Generally, observations represent a state close to reality, but they are spatially sparse. In contrast, fluid model outputs (i.e., background states) include larger errors, but they are spatially dense. DA methods correct these background states with observations in a non-local manner using covariance matrices. For instance, when an observed value exists at a grid point, this observation is used to correct surrounding (i.e., slightly distant) background values that are correlated with the background at that point. Hereafter, the term “non-local” means that an inferred value at a grid point depends on input values in the neighborhood (i.e., at slightly distant grid points). Conversely, “local” means that an inferred value is computed only from the input values at the same point.
@@
The estimation of covariance matrices is also important for DA using autoencoders (including VAEs). In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown [15, 19]. More generally, the background error distributions are non-Gaussian and can be estimated by VAEs [25]. The non-Gaussian distributions learned by VAEs are then incorporated into a traditional DA method [25], namely, the three-dimensional variational (3D-Var [26]) formalism. In the present study, we focus on the non-locality of super-resolution (SR) and utilize SR operators for estimating covariance structures.
@@
SR is originally a technique to enhance the resolution of images. In computer vision, deep learning-based SR has been actively studied [27, 28, 29]. A key characteristic of SR is its non-local nature [30, 31, 32, 33]: neural networks focus on input pixels at low resolution (LR) that are distant from a target pixel at high resolution (HR), but correlated to this target, generating high-quality HR images.
@@
Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA) [34, 35]. One reason for this integration is the increasing availability of HR observations [36, 37, 38]. In typical DA methods, HR observations are assimilated using HR fluid models [39, 40], which require a large amount of computational resources for time integration. In contrast, in SRDA frameworks, background states are computed with LR fluid models. The LR results are then super-resolved to HR using neural networks while assimilating HR observations into these super-resolved backgrounds. SRDA methods are computationally efficient because they do not require numerical integration of HR fluid models.
@@
One advantage of combining SR with DA is that advanced techniques developed in SR can be applied in DA. Numerous SR studies have been conducted not only in computer vision [27, 28, 29] but also in fluid dynamics [41]. In contrast, DA using deep learning has only recently become active [42]. As for SRDA, several research questions remain, such as those regarding its theoretical background. Indeed, previous studies on SRDA [34, 35] did not utilize the non-locality of SR. If theoretical backgrounds for combining SR with DA are clarified, further integration of both will be promoted.
@@
This study proposes a theory of unsupervised SRDA using CVAEs and shows that SR operators can be used to estimate background covariances. This result provides a theoretical foundation for combining SR with DA. We first review an evidence lower bound (ELBO) for DA (Section 2), which serves as the loss function for unsupervised learning. We then develop a new theory of DA using the ELBO and show that this theory is an extension of the 3D-Var (Section 3). Similar to the 3D-Var, our theory requires non-diagonal background error covariance matrices. We further extend our theory by incorporating SR for non-local inference, which eliminates the need to estimate non-diagonal covariance matrices (Section 4). To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section 5). The results are analyzed through comparison with an ensemble Kalman filter (Section 6). Conclusions are presented in Section 7.
2 Review of ELBO for DA
@@
We reformulate an evidence lower bound (ELBO) [43, 44, 45], which is the objective function in the proposed theory. ELBO is also known as the variational lower bound.
@@
The ELBO is derived by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$ [43, 44, 45]:
@@
DA is usually performed via maximum likelihood estimation [1, 2]. Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], we maximize the ELBO, or equivalently minimize $l$ in Eq. (2c) to obtain $q(\bm{z}\mid\bm{x},\bm{y})$. The true state $\bm{z}$ is then inferred from $\bm{x}$ and $\bm{y}$ through $q(\bm{z}\mid\bm{x},\bm{y})$.
3 DA using the ELBO under Gaussianity
@@
3.1 Transformation of the ELBO under Gaussianity
To transform the ELBO, the following Gaussian distributions are introduced:
@@ 3.1 Transformation of the ELBO under Gaussianity
Here, $\bm{a}$ and $V$ denote the vector and matrix given by $\mathsf{A}(\bm{x},\bm{y})$ and $V(\bm{x},\bm{y})$, respectively, $\bm{\varepsilon}$ is an $n_{z}$-dimensional standard normal variable, and $V^{1/2}$ is obtained by Cholesky decomposition. In Eq. (4b), the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trick [8]), where each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ as in Eq. (5). In the actual implementation, the sample size $n_{\text{sample}}$ can be set to 1 if the batch size is sufficiently large [8]. Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
@@
$\mathbb{E}_{q}$ $\left[\ln\frac{q\left(\bm{z}\mid\bm{x},\bm{y}\right)}{p\left(\bm{z}\mid\bm{x}\right)}\right]$
@@
$l_{\text{Gauss}}$ $=$ $\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}$ (7)
$\quad+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$
where constant terms are omitted and the sample size $n_{\rm sample}$ in Eq. (4b) is set to 1.
3.2 Reduction of the ELBO to the 3D-Var objective function
@@
We start by setting the covariance $V(\bm{x},\bm{y})$ to a constant matrix, where each component is quite small. The variable $\bm{\hat{z}}$ in Eq. (5) is then approximated as $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})=\bm{a}$, and the third term in Eq. (7) becomes constant. Consequently, $l_{\text{Gauss}}$ is reduced to the 3D-Var objective function $l_{\text{3D-Var}}$:
@@
where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$. This $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$. The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq. (2c), represents the difference between the observation vector $\bm{y}$ and the analysis $\bm{a}$ via the observation operator $\mathsf{H}$. The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq. (2c), represents the difference between the analysis $\bm{a}$ and the background state $\bm{x}$. By minimizing $l_{\text{3D-Var}}$, $\bm{x}$ and $\bm{y}$ are integrated to estimate the true state as $\bm{a}$. Since $l_{\text{3D-Var}}$ is derived from the ELBO [i.e., $l_{\text{Gauss}}$ in Eq. (7)], the ELBO describes the balance between $\bm{x}$ and $\bm{y}$ in estimating the true state.
The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$ [1].
@@
where $K$ is known as the Kalman gain matrix. These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values. This non-locality typically originates from the non-zero off-diagonal elements of the covariance matrix $B$. If $B$ is diagonal (i.e., all off-diagonal elements are zero), DA becomes local inference, and spatially sparse observations are incorporated into the inference only at their located grid points.
3.3 Implementation of DA using CVAE

### v5

# Unsupervised Super-Resolution Data Assimilation Using Conditional Variational Autoencoders with Estimating Background Covariances via Super-Resolution

@@ Abstract
This study proposes a theory of unsupervised super-resolution data assimilation (SRDA) using conditional variational autoencoders (CVAEs). We derive an evidence lower bound for unsupervised learning, showing that our theory is an extension of a traditional data assimilation (DA) method, namely the three-dimensional variational (3D-Var) formalism. In contrast to 3D-Var, our theory exploits the non-locality of super-resolution (SR) to learn background covariances without explicitly imposing them for assimilating distant observations. For linear SR, SR operators serve as background error covariance matrices,whereas for nonlinear SR, error backpropagation through SR neural networks induces covariance structures in inference. SRDA can naturally be realized with CVAEs because the loss function for CVAEs is generally an evidence lower bound. By incorporating the SR neural network into the CVAE, the encoder estimates the high-resolution (HR) analysis from HR observations and low-resolution forecasts. The decoder acts as the observation operator by reconstructing the HR observations from the estimated HR analysis. The effectiveness of SRDA was evaluated through numerical experiments using an idealized barotropic ocean jet system. Compared to inference with an ensemble Kalman filter, SRDA demonstrated superior accuracy in HR inference. SRDA was also computationally efficient because it does not require HR numerical integration or ensemble calculations. The findings of this study provide a theoretical basis for integrating SR and DA, which will stimulate further research in this direction.
preprint: AIP/123-QED
I Introduction
@@
Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD).Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time. Uncertainties in model parameters may also lead to discrepancies from reality. DA reduces these errors by correcting forecast values or model parameters based on observational data, thereby being essential for accurate predictions, such as numerical weather prediction.
@@
Recent advances in deep learning are impacting various scientific fields and are also found in DA. One reason is that DA and deep learning use backpropagation based on adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al. (2018); Geer (2021) Another reason is the use of the Bayesian approach.Carrassi et al. (2018); Geer (2021); Bocquet et al. (2020) This approach models the true states of fluid systems by probability distributions based on known data, such as observations and fluid model outputs.
@@
A typical neural network using the Bayesian approach is the variational autoencoder (VAE).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) Like standard autoencoders,Dong et al. (2018) a VAE consists of an encoder and a decoder, both of which are implemented by neural networks. Typically, the encoder transforms the input data into a compressed representation in the latent space, while the decoder reconstructs the original data from these latent variables. A major extension of VAEs is the conditional VAE (CVAE), which is based on conditional probability modeling.Kingma et al. (2014); Sohn, Yan, and Lee (2015) This conditioning allows us to model target variables, such as the true states of systems, conditioned on known data, such as observations.
@@
Applications of VAEs in DA can be categorized into two groups. The first group employs VAEs for dimensionality reduction.Liu, Grana, and de Figueiredo (2021); Mohd Razak et al. (2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) In this approach, the encoder maps fluid model states into a low-dimensional latent space. Statistical DA methods, such as Kalman filters, are then applied in this space. Due to the reduced dimensionality, the computational cost for DA is significantly lowered. This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks.Mack et al. (2020); Amendola et al. (2021); Peyron et al. (2021); Cheng et al. (2022); Liu et al. (2022); Cheng et al. (2024) The second group employs VAEs as generators that learn the probability distributions of flow fields.Grooms (2021); Yang and Grooms (2021); Grooms et al. (2023) In this approach, a fluid model state is encoded, and noise is then added in the latent space, creating various potential latent variables. Ensemble members are obtained by decoding these latent variables back into the physical space. This process facilitates fast computation of covariance matrices used in statistical DA methods. A common characteristic in both groups is that VAEs do not directly perform DA; rather, VAEs are used to make statistical DA methods more efficient. We proposes a new method, categorized into the third group, where VAEs directly perform DA.
@@
In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state.Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Generally, observations represent a state close to reality, but they are spatially sparse. In contrast, fluid model outputs (i.e., background states) include larger errors, but they are spatially dense. DA methods correct these background states with observations in a non-local manner using covariance matrices. For instance, when an observed value exists at a grid point, this observation is used to correct surrounding (i.e., slightly distant) background values that are correlated with the background at that point. Hereafter, the term “non-local” means that an inferred value at a grid point depends on input values in the neighborhood (i.e., at slightly distant grid points). Conversely, “local” means that an inferred value is computed only from the input values at the same point.
@@
The estimation of covariance matrices is also important for DA using autoencoders, including VAEs. In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown.Melinc, Boštjan and Zaplotnik, Žiga (2024); Cheng et al. (2022) More generally, the background error distributions are non-Gaussian and can be estimated by VAEs.Xiao et al. (2025) The non-Gaussian distributions learned by VAEs are then incorporated into a traditional DA method,Xiao et al. (2025) namely, the three-dimensional variational (3D-Var)Lorenc (1986) formalism. In the present study, we focus on the non-locality of super-resolution (SR) and utilize SR operators for estimating covariance structures.
@@
SR is originally a technique to enhance the resolution of images. In computer vision, deep learning-based SR has been actively studied.Ha et al. (2019); Anwar, Khan, and Barnes (2020); Lepcha et al. (2023) A key characteristic of SR is its non-local nature:Dong et al. (2014); Gu and Dong (2021); Chen et al. (2023); Zhou et al. (2023) neural networks focus on input pixels at low resolution (LR) that are distant from a target pixel at high resolution (HR), but correlated to this target, generating high-quality HR images.
@@
Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA).Barthélémy et al. (2022); Yasuda and Onishi (2023) One reason for this integration is the increasing availability of HR observations.Imaoka et al. (2010); Durand et al. (2010); Cifelli et al. (2018) In typical DA methods, HR observations are assimilated using HR fluid models,Li, Wang, and Fu (2019); Honda et al. (2022) which require a large amount of computational resources for time integration. In contrast, in SRDA frameworks, background states are computed with LR fluid models. The LR results are then super-resolved to HR using neural networks while assimilating HR observations into these super-resolved backgrounds. SRDA methods are computationally efficient because they do not require numerical integration of HR fluid models.
@@
One advantage of combining SR with DA is that advanced techniques developed in SR can be applied in DA. Numerous SR studies have been conducted not only in computer visionHa et al. (2019); Anwar, Khan, and Barnes (2020); Lepcha et al. (2023) but also in fluid dynamics.Fukami, Fukagata, and Taira (2023) In contrast, DA using deep learning has only recently become active.Cheng et al. (2023) As for SRDA, several research questions remain, such as those regarding its theoretical background. Indeed, previous studies on SRDABarthélémy et al. (2022); Yasuda and Onishi (2023) did not utilize the non-locality of SR. If theoretical backgrounds for combining SR with DA are clarified, further integration of both will be promoted.
@@
This study proposes a theory of unsupervised SRDA using CVAEs and shows that SR operators can be used to estimate background covariances. This result provides a theoretical foundation for combining SR with DA. We first review an evidence lower bound (ELBO) for DA (Section II), which serves as the loss function for unsupervised learning. We then develop a new theory of DA using the ELBO and show that this theory is an extension of the 3D-Var (Section III). Similar to the 3D-Var, our theory requires non-diagonal background error covariance matrices. We further extend our theory by incorporating SR for non-local inference, which eliminates the need to estimate non-diagonal covariance matrices (Section IV). To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section V). The results are analyzed through comparison with an ensemble Kalman filter (Section VI). Conclusions are presented in Section VII.
II Review of ELBO for DA
@@
We reformulate an evidence lower bound (ELBO),Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017) which is the objective function in the proposed theory. ELBO is also known as the variational lower bound.
@@
The ELBO is derived by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$:Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017)
@@
DA is usually performed via maximum likelihood estimation.Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], we maximize the ELBO, or equivalently minimize $l$ in Eq. (2c) to obtain $q(\bm{z}\mid\bm{x},\bm{y})$. The true state $\bm{z}$ is then inferred from $\bm{x}$ and $\bm{y}$ through $q(\bm{z}\mid\bm{x},\bm{y})$.
III DA using the ELBO under Gaussianity
@@
III.1 Transformation of the ELBO under Gaussianity
We make two assumptions to derive the loss function for DA: (A) the ELBO in Eq. (2c) and (B) the following Gaussian distributions:
@@ III.1 Transformation of the ELBO under Gaussianity
Here, $\bm{a}$ and $V$ denote the vector and matrix given by $\mathsf{A}(\bm{x},\bm{y})$ and $V(\bm{x},\bm{y})$, respectively, $\bm{\varepsilon}$ is an $n_{z}$-dimensional standard normal variable, and $V^{1/2}$ is obtained by Cholesky decomposition. In Eq. (4b), the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trick),Kingma and Welling (2014) where each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ as in Eq. (5). In the actual implementation, the sample size $n_{\text{sample}}$ can be set to 1 if the batch size is sufficiently large.Kingma and Welling (2014) Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
@@
$l_{\text{Gauss}}$ $=$ $\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$ (7)
where constant terms are omitted and the sample size $n_{\text{sample}}$ in Eq. (4b) is set to 1.
III.2 Reduction of the ELBO to the 3D-Var objective function
@@
To show this fact, we make further one assumption: the covariance $V(\bm{x},\bm{y})$ is a constant matrix, where each component is quite small. The variable $\bm{\hat{z}}$ in Eq. (5) is then approximated as $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})=\bm{a}$, and the third term in Eq. (7) becomes constant. Consequently, $l_{\text{Gauss}}$ is reduced to the 3D-Var objective function $l_{\text{3D-Var}}$:
@@
where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$ [i.e., $\bm{a}=\mathsf{A}(\bm{x},\bm{y})$]. This $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$. The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq. (2c), represents the difference between the observation vector $\bm{y}$ and the analysis $\bm{a}$ via the observation operator $\mathsf{H}$. The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq. (2c), represents the difference between the analysis $\bm{a}$ and the background state $\bm{x}$. By minimizing $l_{\text{3D-Var}}$, $\bm{x}$ and $\bm{y}$ are integrated to estimate the true state as $\bm{a}$. Since $l_{\text{3D-Var}}$ is derived from the ELBO [i.e., $l_{\text{Gauss}}$ in Eq. (7)], the ELBO describes the balance between $\bm{x}$ and $\bm{y}$ in estimating the true state.
The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$:
@@
where $K$ is known as the Kalman gain matrix.Asch, Bocquet, and Nodet (2016) These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values. This non-locality typically originates from the non-zero off-diagonal elements of the covariance matrix $B$. If $B$ is diagonal (i.e., all off-diagonal elements are zero), DA becomes local inference, and spatially sparse observations are incorporated into the inference only at their located grid points.
III.3 Implementation of DA using CVAE